can you explain me from basics for problems with conductors having cavities and cavities containing charges and to find out 1.electric field at different points 2.to know about the uniformity of the charges induced on different surfaces of conductor and cavities.can u give detailed discussion about this and a problem which contains all these concepts?

Grade:12

2 Answers

Aman Bansal
12 years ago

Dear Shna,

1. The problem statement, all variables and given/known data

A point charge q1 = -6.00nC is at the origin, and a second point charge q2 = +4.00nC is on the x-axis at x = 0.800 m. Find the electric field at each of the following points on the x-axis; a.) 0.200m b.) x = 1.20m c.) x = -0.200 m

2. Relevant equations

E = kq1q2/r^2


3. The attempt at a solution

i did a little research and saw this:

E = E1 + E2

so i did it. E1 = (9E9)(-6.00E-9)/(0.2)2 + E2 = (9E9)(+4.00E-9)/(0.6)2

= -1.350x103 + 1.00X102 = -1.25x103

Best Of luck

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

R shna
12 years ago

your solutuion doesnt answer my question